OpenStudy (anonymous):

Which property justifies the following equation? 7(6 + 0 + 5) = 7(6 + 5) A. distributive B. commutative C. associative D. identity

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

i think its A but im not sure

OpenStudy (skullpatrol):

Zero is called the identity element for addition because the sum of 0 and any number is identical to that number. a + 0 = 0 + a = a

OpenStudy (anonymous):

oooo thabk you so much

OpenStudy (skullpatrol):

Thanks for asking :)

OpenStudy (anonymous):

can you help me with another question

OpenStudy (anonymous):

Which property justifies the following equation? 7[6+5+(-6)]=7[6+(-6)+5] A. distributive B. commutative C. associative D. identity

OpenStudy (skullpatrol):